OpenStudy (anonymous):

Compute the monthly payments for an add-on interest loan of $1,200, with an annual interest rate of 9 percent and a term of 2 years. Round to the nearest cent as needed.

OpenStudy (campbell_st):

find the interest using the simple interest formula I = Prn P = 1200 r = 9/100 and n = 2 I = 216 the interest on the loan is $216 the total to be repaid is 1200 + 216 Total = 1416 the monthly repayment is 1416/24 since it is 2 years or 24 months 1416/24 = 59 The monthly repayment on the loan is $59.00

OpenStudy (anonymous):

Thank you!!!!

OpenStudy (dumbcow):

nevermind just looked up what an add-on loan is ...it uses simple interest

OpenStudy (campbell_st):

lol... like me... I'm use to simple interest